OpenStudy (anonymous):

Given the equation y − 4 =3/4 (x + 8) in point-slope form, identify the equation of the same line in slope-intercept form. −3/4x + y = 10 3x − 4y = −40 y = 3/4x + 12 y = 3/4x + 10

OpenStudy (anonymous):

Y is never a number, M is the slope, X is never a number, and B is the Yintercept you can eliminate A and B because they are the wrong format now check for the Yintercept, by plugging 0 in for the X in the given equation (because the Yintercept is when X=0) y-4=3/4(0+8) solve for Y y-4=3/4(8) y-4=9 y=10

OpenStudy (anonymous):

y-4=3/4x+6 y=3/4x+6+4 y=3/4x+10